Fermi National Laboratory
 

Accelerator Update for November 12, 2007 - November 14, 2007

 

Monday November 12

The day shift began with Operations monitoring the Tevatron (TeV) store 5714 and the Recycler stash (182.0E10), with the Antiproton Source (Pbar) stacking (87.0E10), with the Main Injector (MI), Booster, and Linac operating normally, with the Collider Detector at Fermilab (CDF) and D-Zero (D0) using the colliding beams, and with the Meson MTest, and SciBooNE / MiniBooNE experiments taking beam.

Operations terminated store 5714 at 8:08 AM.

Operations reported at 9:30 AM that cryo system technicians had completed the lead flow orifice work.

TeV experts began a helix study and measurement around 10 AM.

Operations reported at 10:29 AM that the nitrogen delivery had arrived and the truck was filling the TeV sector A0 nitrogen DEWAR.

The Neutron Therapy Facility reported at 10:51 AM that they had treated their last patient for the day.

Operations prepared Pbar at 10:58 AM for the first of several accesses to AP0 for the T972 experiment. (The purpose of this experiment is to measure the effectiveness of shielding.)

TeV studies ended at 11:50 AM.

At 12:19 PM, the 750 KeV area buncher tripped off. It reset.

Operations put the TeV into shot setup at 12:20 PM.

Operations established store 5716 at 2:34 PM with an initial luminosity of 84.5E30.

At 2:39 PM, a separator spark at TeV sector B11 caused a 50% loss of luminosity.

At 2:57 PM, CDF reported that they had to turn off their detector due to problems with the antiproton beam. An expert repositioned the collimators and CDF was back on line 15 minutes later.

A MI expert began a slipstacking study at 2:58 PM.

T972 experimenters reported at 4:05 PM that they were done for the day.

The MI study ended at 4:24 PM.

The MTest experimenters reported around 6:40 PM that they and their experiment had completed their data taking and no longer needed the beam. The next experiment to use MTest will begin on December 10, 2007.

At 6:56 PM, three Linac RF stations (LRF2, 3, & 4) tripped off on a switch tube indicator. They all reset.

At 7:13 PM, the Linac buncher tripped off. It reset.

At 8:05 PM, a MI RF station (MIRF10) tripped off for the first of three times in five minutes. It reset each time.

A TeV expert completed a helix change at 10:03 PM to increase luminosity. Operations adjusted collimators at CDF’s request.

Operations reported at midnight that the Recycler had stashed antiprotons three times during the evening shift.

Tuesday November 13

The midnight shift began with Operations monitoring store 5716 and the stash (203.8E10), with Pbar stacking (42.4E10), and with Sci /MiniBooNE taking beam.

At 12:22 AM, the Linac buncher tripped off. It reset.

At 2:42 AM, a MI crate lost its connection. Operators investigated and found that its power supply had failed. Operators replaced the power supply.

At 4:21 AM, the 750 KeV area buncher tripped off. It reset.

At 5:09 AM, LRF3 and 5 tripped off. Operators reset the stations, but LRF5 tripped off several more times. Eventually everything reset.

At 5:20 AM, the Tevatron Electron Lens (TEL1) tripped off on a vacuum indication. It reset.

At 7:43 AM, Operations reported a resistivity alarm for the #3 CUB column.

At 7:56 AM, the 750 KeV area buncher tripped off. It reset.

At 8:23 AM, store 5716 suffered a sector D1 quench. The store had been scheduled for termination at 8:30 AM.

Operations prepared CDF for access at 9:05 AM.

At 9:11 AM, a power supply (E:MBEXPS) tripped off Sci/MiniBooNE. This happened twice in 17 minutes. An expert investigated and made some adjustments.

CDF came out of access at 9:45 AM.

Operations prepared Pbar at 10:22 AM for the first of several accesses throughout the day to AP0 by the T972 experimenters.

A Linac expert turned off LRF5 at 10:34 AM to install some diagnostic equipment. The station was back online within five minutes.

Booster experts began tuning their machine at 10:54 AM.

A MI expert began a one-hour slipstacking study at 11:05 AM.

SwitchYard (SY) experts began a study at 11:08 AM.

Operations put the TeV through a dry squeeze at 11:52 AM.

TeV experts began a study at 1:03 PM.

SY studies ended at 3 PM.

An expert reported at 3:15 PM that the TeV sector A0 nitrogen DEWAR had been filled and that for now it would be filled once every day.

TeV studies ended at 4:04 PM.

Operations prepared the CDF and D0 collision halls for access at 4:12 PM.

TeV experts put all the TeV separators into conditioning mode at 4:29 PM.

D0 and CDF were both out of access by 6:34 PM.

Operations put the TeV through a dry squeeze at 6:56 PM.

At 8:51 PM, Operations reported problems with stacking. An expert investigated.

Wednesday November 14

The midnight shift began with Operations monitoring the stash (199.4E10), with Pbar stacking (249.0E10), and with MTest and Sci /MiniBooNE taking beam.

At 1:11 AM, the 750 KeV area buncher tripped off. It reset.

Operations prepared TeV and D0 for an access at 2:41 AM.

At 3:23 AM, the 750 KeV area buncher tripped off. It reset.

At 4:44 AM, the 750 KeV area buncher tripped off. It reset.

At 6:16 AM, the 750 KeV area buncher tripped off. It reset.

The Pbar stack reached 258.48 milliamps (mA) at 7:56 AM, which was a record stack. The previous stack record was 245mA.

The Plans for Wednesday

The plans for today are to allow the TeV separator conditioning to continue. The Run Coordinator would like to see a store in the TeV later this afternoon or early this evening. There may be a 12-hour access tomorrow (11-15-07) into MI to repair LCW leaks on the bus around the new MI collimators. The Run Coordinator won’t decide about the access until later today.


Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 11/14/2007 -->    email Fermilab